OpenStudy (anonymous):

Right triangle AC= 16m AB= ?m

OpenStudy (anonymous):



OpenStudy (anonymous):

its hard to draw pictures on here

OpenStudy (anonymous):

2 AC^2 = 16^2 AC = sqrt 16^2 / 2 AC = sqrt 128 = 11.31 correct to 2 D P's i know! ( about drawing pics)

OpenStudy (anonymous):

\[AC=\sqrt{128}=8\sqrt{2}\]

OpenStudy (anonymous):

huh? well there is 3 answers I can pick from 8^2 or 16^2 or neither